OpenStudy (anonymous):

Help! Need answer asap!!Stacey's goal is to workout for 200 minutes a week. She has been increasing her time a little each week. Her times for the first four are week 1: 70 Week2: 80 Week3: 95 Week4: 115 In what week will she reach her goal? Explain your thinking at each step and solution.

OpenStudy (anonymous):

She will reach her goal during week 7 with 205 minutes working out. I got this by figuring out the pattern- Week 1 to week 2: +10 Week 2 to week 3: +15 Week 3 to week 4: +20 So, I came to realize that it was adding 5, so I just added until I got past 200 minutes. 115+25= 140+30= 170+ 35= 205!!!

OpenStudy (anonymous):

Wow thank you!